# Joint Sampling Frequency Offset Estimation and Compensation Based on the Farrow Structure

### v2

@@ I Introduction
In digital communication systems, accurate synchronization is crucial for the correct reception and interpretation of signals. Among various synchronization challenges, sampling frequency offset (SFO) critically impacts the system performance as it leads to inter-carrier and inter-symbol interference [1, 2]. Most of the existing methods for SFO estimation are designed for orthogonal frequency division multiplexing (OFDM) signals, and therefore carry out the estimation in the frequency domain [3, 4, 5, 6, 7, 8, 9]. Several methods consider joint estimation of SFO and carrier frequency offset (CFO) due to their interrelated effects [5, 6, 7, 8, 9]. Alternatively, time-domain SFO estimation methods have been proposed to reduce the implementation complexity [10, 11], but their dependency on other errors has not been investigated in these studies. While these techniques have demonstrated efficacy for OFDM based systems, they can require substantial computational resources for estimation in systems that do not rely on the fast Fourier transform (FFT). Additionally, many SFO estimation methods assume perfect synchronization of the first samples and, therefore, do not account for sampling time offset (STO), which, in conjunction with SFO, introduces an additional challenge for accurate estimation. Apart from this, existing techniques typically work on a complex-valued signal, consequently increasing the number of real multiplications111 One complex multiplication requires a minimum of three real multiplications [12]. required for the estimation.
@@ III Proposed Time-Domain SFO Estimation Based on the Farrow Structure and Newton Method
$\mathbf{H}=\begin{bmatrix}\displaystyle\frac{\partial^{2}F}{\partial\Delta^{2}}&\displaystyle\frac{\partial^{2}F}{\partial\Delta\partial\varepsilon}\\ \\ \displaystyle\frac{\partial^{2}F}{\partial\varepsilon\partial\Delta}&\displaystyle\frac{\partial^{2}F}{\partial\varepsilon^{2}}\end{bmatrix}$ $=\begin{bmatrix}\displaystyle\sum_{n}n^{2}\frac{\partial^{2}F_{n}}{\partial d^{2}}&\displaystyle\sum_{n}n\frac{\partial^{2}F_{n}}{\partial d^{2}}\\ \\ \displaystyle\sum_{n}n\frac{\partial^{2}F_{n}}{\partial d^{2}}&\displaystyle\sum_{n}\frac{\partial^{2}F_{n}}{\partial d^{2}}\end{bmatrix},$ (7)
$\mathbf{g}=\begin{bmatrix}\displaystyle\frac{\partial F}{\partial\Delta}&\displaystyle\frac{\partial F}{\partial\varepsilon}\end{bmatrix}^{\top}$ $=\begin{bmatrix}\displaystyle\sum_{n}n\frac{\partial F_{n}}{\partial d}&\displaystyle\sum_{n}\frac{\partial F_{n}}{\partial d}\end{bmatrix}^{\top},$ (8)
@@
$\frac{\partial F_{n}}{\partial d}$ $=\Bigg{(}\sum_{k=0}^{L}d^{k}u_{k}(n)-x_{0}(n)\Bigg{)}\Bigg{(}\sum_{k=1}^{L}kd^{k-1}u_{k}(n)\Bigg{)},$ (9)
@@
×$\times\Bigg{(}\sum_{k=2}^{L}k(k-1)d^{k-2}u_{k}(n)\Bigg{)}.$ (10)
@@ IV-B Implementation Complexity of the Proposed Estimator
The implementation complexity of the proposed estimator comprises the update computations as described in (6), with the subsequent matrix inversion, and the computation of the derivatives to obtain the elements in (7)–(8), via the proposed method for the sum of multiplications using cascaded accumulators outlined in Section IV-A. To assess the complexity, we distinguish between general multiplications and constant multiplications since the latter can be efficiently implemented using adders and shift operators. Additionally, some parts of the estimation algorithm, specifically the subfilter outputs $u_{k}(n)$, $k=0,\dots,L$, multiplications by $d$, and additions required to compute the output $y_{c}(n,d)$, are utilized for compensation (in Fig. 2, this part is highlighted with a gray rectangle). Given that they are inherently required for this purpose, their cost is excluded from the overall complexity of the estimator. Therefore, considering that every accumulator in (13)–(14) needs $N-1$ additions, the proposed estimation algorithm requires $(L+2)\times N+8$ general multiplications, $(2L-3)\times N+5$ real multiplications, $(2L+6)\times N+3$ additions, and one division per batch of $N$ samples. This includes the cost of the update in (6), which totally requires eight general multiplications, five additions and one division. Further in regular Farrow structures, $L$ is a small number, typically between one and five.
@@ V Results
In this section, we evaluate the performance of the proposed method by applying it to different types of bandlimited signals. The goal is to synchronize two signals sampled at $f_{1}$ and $f_{0}$, where $f_{0}$ serves as a reference with a normalized sampling frequency of one. Examples 1–3 demonstrate specific instances from evaluation across 1 00010001\,0001 000 signals for each case. For the SFO and STO set as $\Delta=-200$ ppm and $\varepsilon=0.03$, respectively, the estimation error is upper bounded by a $3$% error margin around the target values, with the majority ($\approx 90$%) of the cases achieving estimation errors upper bounded by $1$%. The signal-to-noise ratio (SNR) is $60$ dB, and $N=256$ samples are used for the estimation.
@@
Example 3: To illustrate the effectiveness of the proposed method in the presence of other errors, $5\%$ CFO and phase offset (PO) are introduced to an OFDM signal with 1 53615361\,5361 536 out of 2 04820482\,0482 048 subcarriers, and 16-QAM. Here, the estimation of SFO and STO also requires one iteration to obtain $\widehat{\Delta}=-201.03$ ppm and $\widehat{\varepsilon}=0.0298$, by using only the real component of the signal for the SFO estimation. Figure 4 shows the constellation diagrams before, after SFO compensation, and after subsequent CFO and PO compensation.
Example 4: Further, Fig. 5 quantifies the performance of the estimator by plotting the normalized mean-squared error (NMSE) and bit error rate (BER) achieved versus SNR for 1 00010001\,0001 000 OFDM signals with ${\Delta}=-200$ ppm and ${\varepsilon}=300$ ppm per SNR value, demonstrating the estimator’s effectiveness even in presence of low SNRs. As in previous examples, one iteration is sufficient.
@@
Example 5: Lastly, the performance of the proposed estimator within a wide range of SFOs is analyzed. Here, OFDM signals using 16-QAM are generated with fixed $\varepsilon=300$ ppm and varying $\Delta$. Figure 6 shows the NMSE versus $\Delta$ across 1 00010001\,0001 000 signals per $\Delta$ value, demonstrating the estimator’s performance under different SFOs. It is seen that the compensated signals are close to the reference. Here, the NMSE is bounded by an SNR of $30$ dB.

### v3

@@ I Introduction
In digital communication systems, accurate synchronization is crucial for the correct reception and interpretation of signals. Among various synchronization challenges, sampling frequency offset (SFO) critically impacts the system performance as it leads to inter-carrier and inter-symbol interference [1, 2]. Most of the existing methods for SFO estimation are designed for orthogonal frequency division multiplexing (OFDM) signals, and therefore carry out the estimation in the frequency domain [3, 4, 5, 6, 7, 8, 9]. Several methods consider joint estimation of SFO and carrier frequency offset (CFO) due to their interrelated effects [5, 6, 7, 8, 9]. Alternatively, time-domain SFO estimation methods have been proposed to reduce the implementation complexity [10, 11], but their dependency on other errors has not been investigated in these studies. While these techniques have demonstrated efficacy for OFDM based systems, they can require substantial computational resources for estimation in systems that do not rely on the fast Fourier transform (FFT). Additionally, many SFO estimation methods assume perfect synchronization of the first samples and, therefore, do not account for sampling time offset (STO), which, in conjunction with SFO, introduces an additional challenge for accurate estimation. Apart from this, existing techniques typically work on a complex-valued signal, consequently increasing the number of real multiplications [^1] One complex multiplication requires a minimum of three real multiplications [12]. required for the estimation.
@@ III Proposed Time-Domain SFO Estimation Based on the Farrow Structure and Newton Method
$\hskip-4.26773pt\mathbf{H}=\begin{bmatrix}\displaystyle\frac{\partial^{2}F}{\partial\Delta^{2}}&\displaystyle\frac{\partial^{2}F}{\partial\Delta\partial\varepsilon}\\ \\ \displaystyle\frac{\partial^{2}F}{\partial\varepsilon\partial\Delta}&\displaystyle\frac{\partial^{2}F}{\partial\varepsilon^{2}}\end{bmatrix}$ $=\begin{bmatrix}\displaystyle\sum_{n}n^{2}\frac{\partial^{2}F_{n}}{\partial d^{2}}&\displaystyle\sum_{n}n\frac{\partial^{2}F_{n}}{\partial d^{2}}\\ \\ \displaystyle\sum_{n}n\frac{\partial^{2}F_{n}}{\partial d^{2}}&\displaystyle\sum_{n}\frac{\partial^{2}F_{n}}{\partial d^{2}}\end{bmatrix}\hskip-3.41418pt$, (7)
$\mathbf{g}=\begin{bmatrix}\displaystyle\frac{\partial F}{\partial\Delta}&\displaystyle\frac{\partial F}{\partial\varepsilon}\end{bmatrix}^{\top}$ $=\begin{bmatrix}\displaystyle\sum_{n}n\frac{\partial F_{n}}{\partial d}&\displaystyle\sum_{n}\frac{\partial F_{n}}{\partial d}\end{bmatrix}^{\top}\hskip-2.84544pt$, (8)
@@
$\frac{\partial F_{n}}{\partial d}$ $=\Bigg(\sum_{k=0}^{L}d^{k}u_{k}(n)-x_{0}(n)\Bigg)\Bigg(\sum_{k=1}^{L}kd^{k-1}u_{k}(n)\Bigg),\hskip-3.0pt$ (9)
@@
$\times\Bigg(\sum_{k=2}^{L}k(k-1)d^{k-2}u_{k}(n)\Bigg).$ (10)
@@ IV-B Implementation Complexity of the Proposed Estimator
The implementation complexity of the proposed estimator comprises the update computations as described in (6), with the subsequent matrix inversion, and the computation of the derivatives to obtain the elements in (7)–(8), via the proposed method for the sum of multiplications using cascaded accumulators outlined in Section IV-A. To assess the complexity, we distinguish between general multiplications and constant multiplications since the latter can be efficiently implemented using adders and shift operators. Additionally, some parts of the estimation algorithm, specifically the subfilter outputs $u_{k}(n)$, $k=0,\dots,L$, multiplications by $d$, and additions required to compute the output $y_{c}(n,d)$, are utilized for compensation (in Fig. 2, this part is highlighted with a gray rectangle). Given that they are inherently required for this purpose, their cost is excluded from the overall complexity of the estimator. Therefore, considering that every accumulator in (13)–(14) needs N1N-1 additions, the proposed estimation algorithm requires $(L+2)\times N+8$ general multiplications, $(2L-3)\times N+5$ real multiplications, $(2L+6)\times N+3$ additions, and one division per batch of $N$ samples. This includes the cost of the update in (6), which totally requires eight general multiplications, five additions and one division. Further in regular Farrow structures, $L$ is a small number, typically between one and five.
@@ V Results
In this section, we evaluate the performance of the proposed method by applying it to different types of bandlimited signals. The goal is to synchronize two signals sampled at $f_{1}$ and $f_{0}$, where $f_{0}$ serves as a reference with a normalized sampling frequency of one. Examples 1–3 demonstrate specific instances from evaluation across 1 0001\,000 signals for each case. For the SFO and STO set as $\Delta=-200$ ppm and $\varepsilon=0.03$, respectively, the estimation error is upper bounded by a 33% error margin around the target values, with the majority ($\approx 90$%) of the cases achieving estimation errors upper bounded by 11%. The signal-to-noise ratio (SNR) is $60$ dB, and $N=256$ samples are used for the estimation.
@@
Example 3: To illustrate the effectiveness of the proposed method in the presence of other errors, $5\%$ CFO and phase offset (PO) are introduced to an OFDM signal with 1 5361\,536 out of 2 0482\,048 subcarriers, and 16-QAM. Here, the estimation of SFO and STO also requires one iteration to obtain $\widehat{\Delta}=-201.03$ ppm and $\widehat{\varepsilon}=0.0298$, by using only the real component of the signal for the SFO estimation. Figure 4 shows the constellation diagrams before, after SFO compensation, and after subsequent CFO and PO compensation.
Example 4: Further, Fig. 5 quantifies the performance of the estimator by plotting the normalized mean-squared error (NMSE) and bit error rate (BER) achieved versus SNR for 1 0001\,000 OFDM signals with ${\Delta}=-200$ ppm and ${\varepsilon}=300$ ppm per SNR value, demonstrating the estimator’s effectiveness even in presence of low SNRs. As in previous examples, one iteration is sufficient.
@@
Example 5: Lastly, the performance of the proposed estimator within a wide range of SFOs is analyzed. Here, OFDM signals using 16-QAM are generated with fixed $\varepsilon=300$ ppm and varying $\Delta$. Figure 6 shows the NMSE versus $\Delta$ across 1 0001\,000 signals per $\Delta$ value, demonstrating the estimator’s performance under different SFOs. It is seen that the compensated signals are close to the reference. Here, the NMSE is bounded by an SNR of $30$ dB.